Mathematics 134 – Intensive Calculus for Science 2 Quiz 3 – February 10, 2006

Your Name:

Do all work on this sheet (front and back if needed). You have 15 minutes. No calculators on this quiz.

A) (10) If F (x) = ∫_{0}^{x^{2}} e^{sin(t)} dt, what is F′(x)?

B) (10) Compute ∫ e^{cos(πx)} sin(πx) dx.

C) (10) Compute ∫ \frac{x−3}{(x^{2}−6x+12)^{3}} dx.